OpenStudy (anonymous):

I need some help with Uk, i think i understand the questions. anybody have time?

OpenStudy (anonymous):

Fk=u*k*n*d, i'm questionable on the d, I'm working with a downhill skier problem. We are doing energy and work currently in our chapter.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry I mean i'm using formula Fk=uk*n*d

OpenStudy (anonymous):

d here is the hypotenuse of the triangle, since it is the true distance the force acted. In this case, I'd say you want to find potential energy at the top, and then convert to kinetic energy. What is the question exactly?

OpenStudy (anonymous):

ok so if i do it that way. i get Fn=mgh, so 60*10*132.26=79356

OpenStudy (anonymous):

hold on. Force normal is mgcos theta, or in this case 60*9.81*cos 41.4. The forces acting on the skier in the direction of the ramp are mgsin(theta) down the ramp and the force of friction backwards. The total force acting on the body is the difference of down force minus friction. The total force acts over the distance d mentioned earlier.

OpenStudy (anonymous):

yea ok, i see my issue. let me try again.

OpenStudy (anonymous):

so 450.1 N, then that times uk of .2, 450.1x.2=90.02

OpenStudy (anonymous):

Fk=uk*n so set up properly, .2 *450.1, i only need to consider d of hypotenuse if it asks for velocity at some point, would that be correct?

OpenStudy (anonymous):

If you're trying to find the work done on the skier, then yes. That's all you need to know

OpenStudy (anonymous):

Great thanks for bringing me along.

OpenStudy (anonymous):

no problem